can you help me with this question please? thanks
(a) 12 one rupee coins are distributed at random among 5 beggars A, B, C, D and E. Find the probability that: (1) They get
Answer #1

To calculate the probabilities, we need to calculate the total number of ways in which 12 coins can be distributed among the 5 beggars.

Following formula can be used to calculate the number of ways:

Number of ways=1+-C1

Where,

n = number of coins

r = number of beggars

16x15x14x13 Total number of ways = 12+5-1C34 = 18C4 =- = 1820 ways 4x 3x2x1

A.

We need to find the probability that beggars get 4,2,0,5 and 1 coins respectively. It is clear that distributing coins like this is one of the ways of distributing coins.

Therefore, the probability of distributing 4.2.0,5,1 coins respectively is 1820

B.

We need to distribute 12 coins in such a way that every beggar gets at least 2 coins. After distributing 2 coins to every 5 beggars, there are still 2 coins left.

The number of ways with which the remaining two coins can be distributed among 5 beggars can be calculated as follows:

6x 5x 4x3 Number of ways = 245-1C3-1 = C4 = -= 15 ways 4x3x2x1 The probability that every beggar get atleast 2 coins = - 1820

C.

We need to distribute coins in such a way that no beggar goes empty-handed. After distributing at least 1 coin to each beggar and 7 coins are remaining.

The number of ways with which the remaining 7 coins can be distributed among 5 beggars can be calculated as follows: